# A General Latent Embedding Approach for Modeling Non-uniform High-dimensional Sparse Hypergraphs with Multiplicity

### v1

@@ 4 Simulation studies
In this section, we conduct simulation studies to evaluate the performance of the proposed method and validate our theoretical findings. To compute the estimators, we employ a projected gradient descent algorithm (Ma et al., 2020) and use a universal singular value thresholding strategy (Chatterjee, 2015) for initialization. Details of the algorithms are in the supplementary materials. We consider various simulation setups where the embedding vectors are generated from a mixture of truncated multivariate Gaussian distributions. Specifically, when considering a $K$-dimensional latent space, we first randomly split the $n$ vertices into $K$ disjoint almost equal-sized groups, where “almost” means the sizes can differ by $1$ when $K$ does not divide $n$. Vertex embeddings are then independently generated as follows. For vertex $i$, if it is in the $k$th group, $\mathbf{z}^{*}_{i}$ would be generated from ${\cal N}_{[-1,1]}(\mathbf{e}_{k},\boldsymbol{\Sigma}_{\mathbf{z}})$, where $\mathbf{e}_{k}\in\{0,1\}^{K}$ is a column vector with the $k$th element being $1$ while other entries being 00, $\boldsymbol{\Sigma}_{\mathbf{z}}$ is the covariance matrix of the embedding vectors, and $[-1,1]$ indicates truncation around the mean by deviation 1 on each coordinate. The embedding vectors $\{\mathbf{f}^{*}_{j}\}_{j\in[m]}$ of the hyperlinks are centered i.i.d. realizations of ${\cal N}_{[-1,1]}(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{f}})$ so that $\sum_{j\in[m]}\mathbf{f}^{*}_{j}=\mathbf{0}$. For both $\boldsymbol{\Sigma}_{\mathbf{z}}$ and $\boldsymbol{\Sigma}_{\mathbf{f}}$, we adopt the autoregressive design where $\Sigma_{\mathbf{z},ij}=\Sigma_{\mathbf{f},ij}=0.2*\rho^{|i-j|}$ with $\rho\in\{0,0.5\}$. In particular, $\rho=0$ gives independent dimensions of the embedding vectors. The vertex degree heterogeneity parameters $\{\alpha_{i}^{*}\}\subset\mathbb{R}$ are i.i.d. realizations of $\mathrm{Uniform}[-1,1]$ and the order-adjusting parameter $\beta^{*}_{m,n}\in\mathbb{R}$ will be specified in each simulation setup. The hyperlinks $\mathbf{y}_{j}$’s are generated independently via $\mathbf{y}_{j}\sim\mathrm{multiBernoulli}\big{\{}\sigma(\boldsymbol{\theta}_{j}^{*})\big{\}}$, where $\boldsymbol{\theta}^{*}_{j}=\beta^{*}_{m,n}\mathbf{1}_{n}+\boldsymbol{\alpha}^{*}+\mathbf{Z}^{*}\mathbf{f}^{*}_{j}$. In Section 4.1, we investigate the effect of sample size, the order-adjusting parameter and the dimension of the latent space on the estimation of $\boldsymbol{\Theta}^{*}=(\boldsymbol{\theta}^{*}_{1},\ldots.\boldsymbol{\theta}^{*}_{m})^{\top}\in\mathbb{R}^{m\times n}$. In Section 4.2, we construct asymptotic confidence intervals for the individual parameters $(\boldsymbol{\alpha}^{\dagger},\mathbf{F}^{\mathbf{G}},\mathbf{Z}^{\mathbf{G}})$ and $(\boldsymbol{\Theta}^{*},\mathbf{P}^{*})$ based on the asymptotic normality results in Section 3.2, and evaluate their finite sample performance.
@@ 4.1 $\boldsymbol{\Theta}$ estimation
We further investigate the effect of the order-adjusting parameter under different latent space dimensions. We let $K\in\{2,3,4\}$, $m=10n=5,000$ and set $-\beta^{*}_{m,n}\in\{0.5,1,\cdots,3.5,4\}$. The results are displayed in Figure 3. We can see that as the value of $-\beta^{*}_{m,n}$ gets bigger, the estimation error grows proportional to $\exp(-\beta^{*}_{m,n})$. The effects of the dimension of the latent embedding space $K$ and the correlation structure among various dimensions are similar to those in Figure 2.
@@ 4.2 Confidence intervals
Figure 4: Empirical coverage for $\{\alpha^{\dagger}_{i}\}_{i\in[n]}$, $\{Z_{ik}^{*}\}_{i\in[n],k\in[2]}$,$\{F_{jk}^{*}\}_{j\in[m],k\in[2]}$ and $\{\theta_{ii}^{*}\}_{i\in[m]}$, along with interval lengths. Each value comes from $1,000$ Monte Carlo repetitions.
@@
In this section, we evaluate the finite sample performance of the constructed confidence intervals for our model parameters, including the embedding vectors, the un-centered degree heterogeneity parameters, $\theta_{ji}$’s, and $p_{ji}$’s. We set $K=2$ and vary $m=n\in\{400,800,1500,3000,6000\}$, $\beta^{*}_{m,n}\in\{0,-1\}$. After generating $(\beta^{*}_{m,n},\boldsymbol{\alpha}^{*},\mathbf{Z}^{*},\mathbf{F}^{*})$, $({\boldsymbol{\alpha}}^{\dagger},\mathbf{Z}^{*},\mathbf{F}^{*})$ are reconstructed accordingly. The $95\%$ confidence intervals are constructed following the methodologies outlined in Section 3.2. Figure 4 displays the empirical coverage values and confidence interval lengths for $(\boldsymbol{\alpha}^{\dagger},\mathbf{f}^{*},\mathbf{z}^{*},\boldsymbol{\Theta}^{*})$ and Figure 5 presents the results for $\mathbf{P}^{*}$, with both figures considering a range of $(m,n)$ values. They are presented in two separate plots due to the differing scales of $(\boldsymbol{\alpha}^{\dagger},\mathbf{f}^{*},\mathbf{z}^{*},\boldsymbol{\Theta}^{*})$ and $\mathbf{P}^{*}$, making it more informative to display their interval lengths separately. In Figure 4, the top box plots summarize empirical coverage values for each specified coverage target. For example, the “F” coverage target at $n=3000$ aggregates $6000$ empirical coverage values for $\{F_{jk}^{*}\}_{j\in[3000],k\in[2]}$. For the case where the coverage target is “Theta” or the probability term, the box plots only display empirical coverage values for $\theta_{ji}^{*}$ with $j=i$, to avoid overburdening the plot software with all $(j,i)$ pairs. The empirical coverage values are based on 1000 Monte Carlo simulations. The box plots of the confidence interval lengths are constructed similarly.
We can see that when $\beta^{*}_{m,n}=0$, the empirical coverage values for different coverage targets all approach the nominal $95\%$ level, as $m$ and $n$ increase. When $\beta^{*}_{m,n}=-1$, the empirical coverage values are not as favorable, highlighting the increased difficulty in making inference in the presence of sparsity. Furthermore, as $(m,n)$ increases, the lengths of the confidence intervals notably shrink. For fixed $\beta^{*}_{m,n}$, the length of the intervals appears proportional to $(m\wedge n)^{-1/2}$, which aligns well with our theory.

### v2

@@ 4 Simulation studies
In this section, we conduct simulation studies to evaluate the performance of the proposed method and validate our theoretical findings. To compute the estimators, we employ a projected gradient descent algorithm (Ma et al., 2020) and use a universal singular value thresholding strategy (Chatterjee, 2015) for initialization. Details of the algorithms are in the supplementary materials. We consider various simulation setups where the embedding vectors are generated from a mixture of truncated multivariate Gaussian distributions. Specifically, when considering a $K$-dimensional latent space, we first randomly split the $n$ vertices into $K$ disjoint almost equal-sized groups, where “almost” means the sizes can differ by $1$ when $K$ does not divide $n$. Vertex embeddings are then independently generated as follows. For vertex $i$, if it is in the $k$th group, $\mathbf{z}^{*}_{i}$ would be generated from ${\cal N}_{[-1,1]}(\mathbf{e}_{k},\boldsymbol{\Sigma}_{\mathbf{z}})$, where $\mathbf{e}_{k}\in\{0,1\}^{K}$ is a column vector with the $k$th element being $1$ while other entries being 0, $\boldsymbol{\Sigma}_{\mathbf{z}}$ is the covariance matrix of the embedding vectors, and [1,1][-1,1] indicates truncation around the mean by deviation 1 on each coordinate. The embedding vectors $\{\mathbf{f}^{*}_{j}\}_{j\in[m]}$ of the hyperlinks are centered i.i.d. realizations of ${\cal N}_{[-1,1]}(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{f}})$ so that $\sum_{j\in[m]}\mathbf{f}^{*}_{j}=\mathbf{0}$. For both $\boldsymbol{\Sigma}_{\mathbf{z}}$ and $\boldsymbol{\Sigma}_{\mathbf{f}}$, we adopt the autoregressive design where $\Sigma_{\mathbf{z},ij}=\Sigma_{\mathbf{f},ij}=0.2*\rho^{|i-j|}$ with $\rho\in\{0,0.5\}$. In particular, $\rho=0$ gives independent dimensions of the embedding vectors. The vertex degree heterogeneity parameters $\{\alpha_{i}^{*}\}\subset\mathbb{R}$ are i.i.d. realizations of $\mathrm{Uniform}[-1,1]$ and the order-adjusting parameter $\beta^{*}_{m,n}\in\mathbb{R}$ will be specified in each simulation setup. The hyperlinks $\mathbf{y}_{j}$’s are generated independently via $\mathbf{y}_{j}\sim\mathrm{multiBernoulli}\big\{\sigma(\boldsymbol{\theta}_{j}^{*})\big\}$, where $\boldsymbol{\theta}^{*}_{j}=\beta^{*}_{m,n}\mathbf{1}_{n}+\boldsymbol{\alpha}^{*}+\mathbf{Z}^{*}\mathbf{f}^{*}_{j}$. In Section 4.1, we investigate the effect of sample size, the order-adjusting parameter and the dimension of the latent space on the estimation of $\boldsymbol{\Theta}^{*}=(\boldsymbol{\theta}^{*}_{1},\ldots.\boldsymbol{\theta}^{*}_{m})^{\top}\in\mathbb{R}^{m\times n}$. In Section 4.2, we construct asymptotic confidence intervals for the individual parameters $(\boldsymbol{\alpha}^{\dagger},\mathbf{F}^{\mathbf{G}},\mathbf{Z}^{\mathbf{G}})$ and $(\boldsymbol{\Theta}^{*},\mathbf{P}^{*})$ based on the asymptotic normality results in Section 3.2, and evaluate their finite sample performance.
@@ 4.1 $\boldsymbol{\Theta}$ estimation
We further investigate the effect of the order-adjusting parameter under different latent space dimensions. We let $K\in\{2,3,4\}$, $m=10n=5,000$ and set $\beta^{*}_{m,n}\in\{0.5,1,\cdots,3.5,4\}$. The results are displayed in Figure 3. We can see that as the value of $\beta^{*}_{m,n}$ gets bigger, the estimation error grows proportional to $\exp(-\beta^{*}_{m,n})$. The effects of the dimension of the latent embedding space $K$ and the correlation structure among various dimensions are similar to those in Figure 2.
@@ 4.2 Confidence intervals
Figure 4: Empirical coverage for $\{\alpha^{\dagger}_{i}\}_{i\in[n]}$, $\{Z_{ik}^{*}\}_{i\in[n],k\in[2]}$$\{F_{jk}^{*}\}_{j\in[m],k\in[2]}$ and $\{\theta_{ii}^{*}\}_{i\in[m]}$, along with interval lengths. Each value comes from $1,000$ Monte Carlo repetitions.
@@
In this section, we evaluate the finite sample performance of the constructed confidence intervals for our model parameters, including the embedding vectors, the un-centered degree heterogeneity parameters, $\theta_{ji}$’s, and $p_{ji}$’s. We set $K=2$ and vary $m=n\in\{400,800,1500,3000,6000\}$, $\beta^{*}_{m,n}\in\{0,-1\}$. After generating $(\beta^{*}_{m,n},\boldsymbol{\alpha}^{*},\mathbf{Z}^{*},\mathbf{F}^{*})$, $({\boldsymbol{\alpha}}^{\dagger},\mathbf{Z}^{*},\mathbf{F}^{*})$ are reconstructed accordingly. The $95\%$ confidence intervals are constructed following the methodologies outlined in Section 3.2. Figure 4 displays the empirical coverage values and confidence interval lengths for $(\boldsymbol{\alpha}^{\dagger},\mathbf{f}^{*},\mathbf{z}^{*},\boldsymbol{\Theta}^{*})$ and Figure 5 presents the results for $\mathbf{P}^{*}$, with both figures considering a range of $(m,n)$ values. They are presented in two separate plots due to the differing scales of $(\boldsymbol{\alpha}^{\dagger},\mathbf{f}^{*},\mathbf{z}^{*},\boldsymbol{\Theta}^{*})$ and $\mathbf{P}^{*}$, making it more informative to display their interval lengths separately. In Figure 4, the upper box plots summarize empirical coverage values for each specified coverage target. For example, the “F” coverage target at $n=3000$ aggregates $6000$ empirical coverage values for $\{F_{jk}^{*}\}_{j\in[3000],k\in[2]}$. For the case where the coverage target is “Theta” or the probability term, the box plots only display empirical coverage values for $\theta_{ji}^{*}$ with $j=i$, to avoid overburdening the plot software with all $(j,i)$ pairs. The empirical coverage values are based on 1000 Monte Carlo simulations. The box plots of the confidence interval lengths are constructed similarly.
We can see that under both $\beta^{*}_{m,n}=0$ and $\beta^{*}_{m,n}=-1$, the empirical coverage values for different coverage targets all approach the nominal $95\%$ level as $m$ and $n$ increase. In both cases, the empirical coverage of $\alpha_{i}^{\dagger}$ achieves $95\%$ with relatively small sample sizes, while the empirical coverage of the latent embeddings is initially below $95\%$ and reaches $95\%$ as the sample size increases, highlighting the difficulty of making inference in the latent space. Furthermore, as $(m,n)$ increase, the lengths of the confidence intervals shrink notably. For fixed $\beta^{*}_{m,n}$, the interval length appears proportional to $(m\wedge n)^{-1/2}$. Moreover, the sparser the hypergraph, i.e., with smaller $\beta^{*}_{m,n}$, the wider the confidence intervals. These empirical findings are consistent with our theory.